The enthalpy of fusion for ice is 6007 J/mol at P =1.00 bar and T=273.15K.

a) What is the molar entropy change (delta S) for fusion of ice at T=273.15 K?

b) What is the Gibbs free energy change (delta G) for the fusion of 1.5 mol of ice at 273.15 K?
Answer #1

Solution :-

Given data

Enthalpy of fusion ( H) = 6007 J/mol

pressure P = 1 bar

Temprature T = 273.15 K

The molar entropy change (delta S) = Enthalpy / melting temprature

= 6007/273.15

= 22 J/mol .K

The Gibbs free energy change (delta G) = H- TS

= 6007-(273.15)(22)

= 0